100g of calcium chloride is dissolved in 1L of aqueous solution. Assume solution density of 1.07g/mL. Calculate the molarity, molality, and percent mass of the solution.
1
Expert's answer
2019-10-28T07:49:40-0400

n(CaCl2)=m/M=100/111=0.9mole

molarity=n(CaCl2)/V(solution)=0.9/1=0.9mole/liter=0.9M

molality=n(CaCl2)/m(solvent)=0.9/(1·1000·1.07)=8,4·10^-4 mole/g

Percent mass=m(CaCl2)/m(solution)=100/(1070+100)=0.085=8.5%